# RRSIS: Referring Remote Sensing Image Segmentation

### v1

@@ IV-C Language-guided Cross-scale Enhancement Module
The main motivation for designing the LGCE module is to leverage language as guidance to adaptively enhance multi-scale visual features. To formally describe LGCE, we denote the encoded visual features from four stages of Swin Transformer as , where , and represent the number of channels, height, and width of feature maps from the -th stage. Among them, is the output of the fourth stage with the highest abstraction but the lowest spatial resolution. Language feature vectors encoded by BERT [41] can be denoted as , where , and represent the number of channels and the count of words, respectively.
@@
where is utilized to divide the given data into two subsets. and denote output language features of and , respectively. and represent output visual features with low and high spatial resolutions. Next, we use and as intermediary tokens to perform cross-scale fusion. Two fully connected (FC) layers and are employed to align dimensions of language features with those of visual features. Specifically, is aligned with , and is aligned with . Further, two multi-head self-attention layers and are exploited to fuse features with different spatial resolutions in Transformer layers. The cross-scale fusion can be formulated as
@@
As shown in Fig. 5, and are concatenated and then fed into the subsequent decoding layers. Afterward, we employ a segmentation head to predict segmentation masks. Following LAVT, the feature decoding part exploits the previous output from the PWAM module in a recursive manner. Two convolutional layers with batch normalization [43] and ReLU are used to construct the feature decoder. Ultimately, a binary segmentation map is predicted by a convolution layer.
@@ V-A Dataset and Evaluation Metrics
The RefSegRS dataset consists of 4,420 image-language-label triples, with 2172 triples in the training set, 431 triples in the validation set, and 1817 triples in the test set. The commonly used metrics in referring image segmentation tasks include overall intersection-over-union (oIoU), mean intersection-over-union (mIoU), and precision at threshold values from 0.5 to 0.9. oIoU is calculated by taking the ratio of the total intersection area to the total union area across all test samples. mIoU is obtained by averaging IoU values between predicted masks and ground truths across all test samples. oIoU tends to give higher importance to larger objects while mIoU considers small and large objects equivalently. Precision at different threshold values measures the proportion of test samples that satisfy a specified IoU threshold.
@@ V-B Competing Methods
LSTM-CNN [4]: It consists of an LSTM network for natural language encoding, a CNN backbone for visual feature extraction, and a fully convolutional upsampling module for pixel-wise segmentation.
ConvLSTM [5]: It adopts a CNN to learn visual features and an LSTM to encode referring expressions. Subsequently, spatial coordinates, visual features, and language features are concatenated together as multi-modal features. Then, multiple convolutional LSTM layers are utilized to refine multi-modal features progressively.
CMSA [6]: It utilizes a CNN for visual feature encoding and keeps individual word vectors encoded by an LSTM. Next, a cross-modal self-attention module is designed to effectively capture long-range dependencies between linguistic and visual features.
BRINet [7]: It employs a CNN to encode visual features and an LSTM to extract linguistic features. Then, a bi-directional cross-modal attention module is utilized to learn cross-modal relationships between linguistic and visual features.
LAVT [8]: It uses Swin Transformer to extract visual features. For the encoding of linguistic features, LAVT uses as the language encoder. As for the fusion of multi-modal features, LAVT integrates the input referring expression, into visual features at intermediate levels in a vision Transformer network.
@@ V-C Implementation Details
Transformer-based models, namely LAVT and the proposed method, are implemented using Pytorch. We set the batch size to 8 at training. The Adam optimizer is used, and the initial learning rate is set to 0.00005. In addition, a weight decay of 0.01 is used for model training, and a window size of 12 is set for the Swin Transformer. For both methods, we train them for 40 epochs and select the best-performing checkpoint according to the performance on the validation set.
@@ V-D1 Comparison with CNN-based methods
To dig deep into the reason causing the performance gap, we opt to visually compare segmentation results produced by different methods. As shown in Fig. 6, targets in vehicle-related samples are extremely cluttered and unevenly distributed. For these examples, CNNs are unable to distinguish small objects due to: 1) the lack of spatial details in features of deep CNN layers; 2) their weak relationship modeling ability. Although multi-level visual features are used in CNN-based methods, their performance on hard examples is still not satisfactory.
@@ V-E Ablation Study and Discussion
The first row represents experimental results of LAVT, which is adopted as our baseline. In the second row, we replace linguistic features with a randomly initialized token to remove language guidance. The clear performance drop indicates the importance of language guidance for multi-scale feature fusion in our model. In the third row, we maintain the remaining parts unchanged while eliminating the cross-scale fusion component. The results show that its segmentation performance in Pr@0.6, Pr@0.7, Pr@0.8, and Pr@0.9 decreases significantly compared with the full model. In the fourth row, the fully-fledged model clearly outperforms the baseline method and set a new state-of-the-art performance on the RefSegRS dataset. The ablation studies can demonstrate the effectiveness of the proposed LGCE module.
@@ VI Conclusion
In this paper, we present a new task of RRSIS. Considering the lack of datasets, we create a new dataset named RefSegRS. The RefSegRS dataset is a collection of 4,420 triplets, and each triplet contains an image, a referring expression, and a corresponding segmentation mask. To fully explore this task, we benchmark a variety of existing methods that are originally designed for natural images in computer vision and offer a comprehensive analysis and valuable insights. Our experimental results reveal that the direct adaptation of existing methods demonstrates limited efficacy in detecting small and scattered objects. Considering this, we propose a novel LGCE module to alleviate this problem. The effectiveness of our LGCE module in detecting such objects is verified by both quantitative and qualitative experimental results.

### v2

@@ IV-C Language-guided Cross-scale Enhancement Module
The main motivation for designing the LGCE module is to leverage language as guidance to adaptively enhance multi-scale visual features. To formally describe LGCE, we denote the encoded visual features from four stages of Swin Transformer as , where , and represent the number of channels, height, and width of feature maps from the -th stage. Among them, is the output of the fourth stage with the highest abstraction but the lowest spatial resolution. Language feature vectors encoded by BERT [38] can be denoted as , where and represent the number of channels and the count of words, respectively.
@@
where is utilized to divide the given data into two subsets. and denote output language features of and , respectively. and represent output visual features. Next, we use and as intermediary tokens to perform cross-scale fusion. Two fully connected (FC) layers and are employed to align dimensions of language features with those of visual features. Specifically, is aligned with , and is aligned with . Further, two multi-head self-attention layers and are exploited to fuse features with different spatial resolutions in Transformer layers. The cross-scale fusion can be formulated as
@@
As shown in Fig. 5, and are concatenated and then fed into the subsequent decoding layers. Afterward, we employ a segmentation head to predict masks. Following LAVT, the feature decoding part exploits the previous output from the PWAM module in a recursive manner. Two convolutional layers with batch normalization [61] and ReLU are used to construct the feature decoder. Ultimately, a binary mask is predicted by a convolution layer.
@@
Methods | FLOPs | PARAMs | MACs |
Baseline (LAVT) | 384.27G | 118.85M | 191.91G |
w/o Language Guidance | 401.34G | 167.37M | 200.44G |
w/o Cross-scale Fusion | 399.44G | 160.48M | 199.49G |
Full Model (LGCE) | 401.34G | 167.37M | 200.44G |
@@ V-A Dataset and Evaluation Metrics
The RefSegRS dataset consists of 4,420 image-language-label triplets, with 2172 triplets in the training set, 431 triplets in the validation set, and 1817 triplets in the test set. The commonly used metrics in referring image segmentation tasks include overall intersection-over-union (oIoU), mean intersection-over-union (mIoU), and precision at threshold values from 0.5 to 0.9. oIoU is calculated by taking the ratio of the total intersection area to the total union area across all test samples. mIoU is obtained by averaging IoU values between predicted masks and ground truths across all test samples. oIoU tends to give higher importance to larger objects while mIoU considers small and large objects equivalently. Precision at different threshold values measures the proportion of test samples that satisfy a specified IoU threshold.
@@ V-B Competing Methods
ConvLSTM [23]: It adopts the ResNet-101 [11] to learn visual features and an LSTM to encode referring expressions. Subsequently, spatial coordinates, visual features, and language features are concatenated together as multi-modal features. Then, multiple convolutional LSTM layers are utilized to refine multi-modal features progressively.
CMSA [24]: It utilizes the ResNet-101 for visual feature encoding and keeps individual word vectors. Next, a cross-modal self-attention module is proposed to effectively capture long-range dependencies between words and spatial regions.
BRINet [25]: It employs the ResNet-101 to encode visual features and an LSTM to extract linguistic features. Then, a bi-directional cross-modal attention module is utilized to learn cross-modal relationships between linguistic and visual features.
LAVT [26]: It uses the Swin Transformer [14] to extract visual features. For the encoding of linguistic features, LAVT uses [38] as the language encoder. As for the fusion of multi-modal features, LAVT integrates the input referring expression into visual features at intermediate levels in a vision Transformer network.
@@ V-C Implementation Details
Transformer-based models, namely LAVT [26] and the proposed method, are implemented using Pytorch. We set the batch size to 8 at training. The AdamW optimizer [62] is used, and the initial learning rate is set to 0.00005. In addition, a weight decay of 0.01 is used for model training, and a window size of 12 is set for the Swin Transformer. For both methods, we train them for 40 epochs and select the best-performing checkpoint according to the performance on the validation set.
@@ V-D1 Comparison with CNN-based methods
To clearly demonstrate the effectiveness of the proposed method, we opt to visually compare segmentation results and their corresponding IoU scores generated by different methods. As shown in Fig. 6, targets in vehicle-related samples are extremely cluttered and unevenly distributed. For these examples, CNNs are unable to distinguish small objects due to: 1) the lack of spatial details in features of deep CNN layers; 2) their weak relationship modeling ability. Although multi-level visual features are used in CNN-based methods, their performance on hard examples is still not satisfactory.
@@ V-E Ablation Study and Discussion
The first row represents experimental results of LAVT, which is adopted as our baseline. In the second row, we replace linguistic features with a randomly initialized token to remove language guidance. The clear performance drop indicates the importance of language guidance for multi-scale feature fusion in our model. In the third row, we maintain the remaining parts unchanged while eliminating the cross-scale fusion component. The results show that its segmentation performance in Pr@0.6, Pr@0.7, Pr@0.8, and Pr@0.9 decreases significantly compared with the full model. In the fourth row, the fully-fledged model clearly outperforms the baseline method and sets a new state-of-the-art performance on the RefSegRS dataset. The ablation studies can demonstrate the effectiveness of the proposed LGCE module.
Regarding the computational complexity, we provide details on the floating point operations (FLOPs), the number of parameters (PARAMs), and multiply-accumulate operations (MACs) for both our model variations and the baseline model. As shown in Table IV, our model incurs only a marginal increase in computational cost compared to the baseline method (LAVT), yet demonstrates significantly improved performance. While the model without cross-scale fusion has lower computational costs than the full model, it leads to a performance decline. In addition, the model without language guidance maintains the same network architecture as the full model, resulting in identical computational costs and the number of parameters.
@@ VI Conclusion
Derived from the concept of referring natural image segmentation, we introduce RRSIS as a novel task within the domain of remote sensing. Considering the lack of datasets, we create a new dataset named RefSegRS. Specifically, we crop tiles from the SkyScapes dataset into images of pixels using a sliding window with a stride of 600 pixels. Then, we generate linguistic expressions for each image with predefined templates. After manually filtering out uninformative image-language-label triplets, we ultimately obtain a dataset consisting of 4,420 image-language-label triplets across 285 scenes. To fully explore this task, we benchmark a variety of existing methods that are originally designed for natural images in computer vision and offer a comprehensive analysis and valuable insights. Our experimental results reveal that the direct adaptation of existing methods demonstrates limited efficacy in detecting small and scattered objects. Considering this, we propose a novel LGCE module to alleviate this problem. The effectiveness of our LGCE module in detecting such objects is verified by both quantitative and qualitative experimental results. For some challenging scenes characterized by cluttered and small objects, LGCE exhibits a remarkable enhancement, achieving an IoU improvement of more than 5% in comparison to the baseline, as shown in Fig. 6.